Not sure where to start on this one hence requiring some aid to solve the question. Working on alegbra, linear equations and systems of equations so it could require any of the above.
You are trying to analyze the time each step in the process takes a worker. There are 3 steps in the process: Material acquisition, building and packaging. Building takes 7 minutes longer than 1/3 of the material acquisition time and packaging requires 3.5 minutes less than half the building time. If a worker produces 72 units in 84 hours of work, how long does each step in the process take per unit produced and what is the ratio of time each step per unit produced?
Even and idea of how to set up the equations would be a huge starting point. Thanks in advance.

Answer by Boreal(15235)

material acquisition time=x
building is (x/3)+7
packaging is [(x/3)+7](1/2)-3.5
Their sum is one unit.
Their time is (84/72) hours per unit or (5040/72) minutes per unit or 70 minutes per unit.
x+(x/3)+7+(1/2)[x/3+7]-3.5=70
Looking at packaging, (x/6)+3.5-3.5=(x/6)
Therefore, x + (x/3) +7 +(x/6)=70
(3/2) x=63
x=63(2/3)=42 minutes material acquisition time
building is 21 minutes
packaging is 7 minutes
The ratio is 6:3:1